OpenStudy (anonymous):

I believe I did this right, but my teacher says its wrong. Can you show me where I went wrong?

OpenStudy (anonymous):

hm so what is the actual question?

OpenStudy (anonymous):

\[\frac{ \sin }{ 1-\cos }+\frac{ \sin }{ 1+\cos }=2\csc\]

OpenStudy (anonymous):

for the second step I got \[\frac{ \sin+\cos \times \sin }{ 1-\cos^2 }+\frac{ \sin-\cos \times \sin }{ 1-\cos^2 }\]

OpenStudy (anonymous):

yeah not bad go one

OpenStudy (anonymous):

\[\frac{ 2\sin }{ 1-\cos^2 }\]

OpenStudy (anonymous):

go on

OpenStudy (anonymous):

\[\frac{ 2\sin }{ \sin^2 }\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[\frac{ 2 }{ \sin }\]

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

\[2\csc=2\csc\]

OpenStudy (anonymous):

correct whats wrong in this

OpenStudy (anonymous):

alright my teacher put a note saying: "This one should begin by finding common denominators and combing the fractions. " so I am confused because I did find a common dendominator

OpenStudy (anonymous):

maybe the common denominator is (1-cos2) and u hve done it right..

OpenStudy (anonymous):

alright well thank you. I will ahve to call her.

OpenStudy (anonymous):

maybe when combining denominator..it has to be one u hve put it in both sides the (1-cos2) it has to be a whole (1-cos 2 )

OpenStudy (anonymous):

\[\frac{1}{1+\cos} + \frac{1}{1-\cos} = \frac{1(1-\cos)+1(1+\cos)}{1-\cos2}\]

OpenStudy (anonymous):

something like this?

OpenStudy (anonymous):

yes I made so that she can see where I made the factors similar. on the bottom I put 1-cos(+cos) and the other side I put 1+cos(-cos)

OpenStudy (anonymous):

then u better call her xD

OpenStudy (anonymous):

I will lol thanks for your help though! :)

OpenStudy (anonymous):

wlcome:)